A force of 295 N is exerted on the end of a wrench in order to apply a torque of 48.4 N·m to a bolt head. The point of application of the force is 21.5 cm from the center of the bolt. What angle does the force make with respect to the wrench handle?
Answer #1

\tau = torque = 48.4 Nm

F = force applied = 295 N

r = distance from center = 21.5 cm = 0.215 m

\theta = angle with wrench handle

Torque is given as

\tau = F r Sin\theta

48.4 = 295 (0.215) Sin\theta

\theta = 49.73